PLEASE HELP ASAP: In a family there are two cars. In a given week, the first car gets an average of 30 miles per gallon, and the second car gets 25 miles per gallon. The two cars combined drive a total of 1700 miles in that week, for a total gas consumption of 60 gallons. How many gallons were consumed by each of the two cars that week? Thank you very much
Answer by josmiceli(6797)

combined miles travelled = (1st car mpg x gals used) + (2nd car mpg x gals used)
let x = gallons used by 1st car. (60 - x)= gallons used by 2nd car
1700 = 30x + 25(60 - x)
200 = 5x
x = 40
(60 - x) = 20
check
1700 = 30 x 40 + 25 x 20
1700 = 1200 + 500
1700 = 1700
OK